limit as x tends to 2,(root of 6-x, minus 2) all over (root of 3-x, minus 1)


1
Expert's answer
2021-05-07T10:21:54-0400

"lim_{x\to 2}\frac{\sqrt{6-x}-2}{\sqrt{3-x}-1}\newline
Putting x=2,\newline
=\frac{\sqrt{4}-2}{\sqrt{1}-1}\newline
=\infty"

Thus, the limit of the given function is "\infty" .